# DOS: Directional Object Separation in Text Embeddings for Multi-Object Image Generation

### v3

@@ Appendix D Directional Encoding Across CLIP Text Embedding Types
One of key motivations for our method is the observation that the difference between two CLIP text embeddings encodes directional information, such as $king-man+woman\approx queen$” (hu2024token). However, prior work has only validated this directional encoding when semantic token and EOT embeddings are used together. Since DOS calculates separation vectors independently for different embedding types—the semantic token for each object noun and the EOT/pooled embedding—here we test whether this directional property holds for each type of embedding individually.
@@ Appendix E Training Data Biases Underlying Generation Failures
The representation of the trained models are known to be significantly influenced by the distribution of their training dataset (nguyen2022quality; qin2023class; xu2024demystifying; khalafi2024constrained). Frequent failures in the two problematic scenarios, Dissimilar Background Biases and Many Objects, can be partially explained by this effect. To investigate this, we analyze all captions in LAION 400M (schuhmann2021laion), a dataset whose filtered subset is believed to be used in training T2I models. Specifically, we compare (i) the co-occurrence rate of object pairs with similar vs. dissimilar background biases, and (ii) the proportion of captions containing varying numbers of noun chunks.
@@
Similar
Background Biases
camel, ostrich 20 / 40655 0.049
@@
Dissimilar
Background Biases
camel, penguin 13 / 65264 0.020
@@
In Table 15, we report the co-occurrence rates for two sets of object pairs: one with similar background biases, and the other with dissimilar biases. The co-occurrence rate is defined using the Jaccard index, calculated as the number of captions containing both objects divided by the number of captions containing at least one of them. The results show that object pairs with dissimilar background biases co-occur significantly less frequently than those with similar biases. In Table 16, we examine how the number of noun chunks varies across prompts. Using spaCy (honnibal2020spacy) to count noun chunks per caption, we find that over 63% of captions contain fewer than three noun chunks. Together, these findings provide evidence that the failures of T2I models in these two scenarios are, at least in part, attributable to biases in the training data distribution.
Table 16: Distribution of noun chunk counts in LAION 400M captions. The proportion of captions with varying numbers of noun chunks, extracted using spaCy (honnibal2020spacy), shows that over 63% of captions contain fewer than three noun chunks, suggesting limited multi-object descriptions in the training data.

### v4

@@ Appendix D Directional Encoding Across CLIP Text Embedding Types
One of key motivations for our method is the observation that the difference between two CLIP text embeddings encodes directional information, such as $king-man+woman\approx queen$” (Hu et al. 2024a). However, prior work has only validated this directional encoding when semantic token and EOT embeddings are used together. Since DOS calculates separation vectors independently for different embedding types—the semantic token for each object noun and the EOT/pooled embedding—here we test whether this directional property holds for each type of embedding individually.
@@ Appendix E Training Data Biases Underlying Generation Failures
The representation of the trained models are known to be significantly influenced by the distribution of their training dataset (Nguyen et al. 2022; Qin et al. 2023; Xu et al. 2024; Khalafi et al. 2024). Frequent failures in the two problematic scenarios, Dissimilar Background Biases and Many Objects, can be partially explained by this effect. To investigate this, we analyze all captions in LAION 400M (Schuhmann et al. 2021), a dataset whose filtered subset is believed to be used in training T2I models. Specifically, we compare (i) the co-occurrence rate of object pairs with similar vs. dissimilar background biases, and (ii) the proportion of captions containing varying numbers of noun chunks.
@@
Similar Background Biases camel, ostrich 20 / 40655 0.049
@@
Dissimilar Background Biases camel, penguin 13 / 65264 0.020
@@
In Table 15, we report the co-occurrence rates for two sets of object pairs: one with similar background biases, and the other with dissimilar biases. The co-occurrence rate is defined using the Jaccard index, calculated as the number of captions containing both objects divided by the number of captions containing at least one of them. The results show that object pairs with dissimilar background biases co-occur significantly less frequently than those with similar biases. In Table 16, we examine how the number of noun chunks varies across prompts. Using spaCy (Honnibal et al. 2020) to count noun chunks per caption, we find that over 63% of captions contain fewer than three noun chunks. Together, these findings provide evidence that the failures of T2I models in these two scenarios are, at least in part, attributable to biases in the training data distribution.
@@
Table 16: Distribution of noun chunk counts in LAION 400M captions. The proportion of captions with varying numbers of noun chunks, extracted using spaCy (Honnibal et al. 2020), shows that over 63% of captions contain fewer than three noun chunks, suggesting limited multi-object descriptions in the training data.